OpenStudy (angelwings996):

Calculate the number of grams of CH_3COOH in the vinegar. The molarity of vinegar that I have gotten is 1.61 M.

OpenStudy (mertsj):

What is the volume of the vinegar that you have gotten?

OpenStudy (angelwings996):

For the volume I got 9.86 ml

OpenStudy (mertsj):

You don't understand. You want to answer the question "how many grams of acetic acid are in the vinegar?" If the solution is 1.61 M it contains 1.61 moles in 1 L of solution. Do you have a whole liter? Is that what you are asking? How many grams of acetic acid in 1 L of a 1.61 M solution?

OpenStudy (angelwings996):

Well, I'm doing a cabbage juice titration lab and one of the questions it is asking is this one but I don't understand how to find the answer.

OpenStudy (mertsj):

If so, then find the molecular weight of acetic acid and multiply y 1.61

OpenStudy (mertsj):

Did you already do the titration?

OpenStudy (angelwings996):

Yes I did

OpenStudy (mertsj):

How many milliliters of 1.61 M acetic acid did you use?

OpenStudy (angelwings996):

Well, the way that I found the molarity of vinegar was: Ma + 9.86ml = 1.14M + 14ml Ma + .00986L = 1.14M + .014L Ma=1.61M

OpenStudy (angelwings996):

@Mertsj

OpenStudy (mertsj):

So are you saying you used 9.86 ml of 1.61 M acetic acid to do the titration?

OpenStudy (angelwings996):

Yes

OpenStudy (mertsj):

1000 ml = 1 L = 1.61 moles acetic acid = 48 g acetic acid 9.86 ml = 9.86/1000 x 1.61 moles acetic acid = .0158746 moles acetic acid = .76 g acetic acid